OpenStudy (anonymous):

How many units apart are the points (5, 20) and (10, 8)?

jimthompson5910 (jim_thompson5910):

d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((10-5)^2+( 8- 20)^2) d = sqrt((5)^2+(-12)^2) d = sqrt(25+144) d = sqrt(169) d = 13 So exact distance between the two points is 13 units.

OpenStudy (anonymous):

wow thanks!

jimthompson5910 (jim_thompson5910):

sure thing